OpenStudy (anonymous):

z^2/Z^2-64=-8 evaluate for the given value of the variable

OpenStudy (anonymous):

How do I cross muliply

OpenStudy (anonymous):

Is the left hand side: \[\frac{z^2}{z^2-64}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Multiply both sides of the question by \(z^2-64\) to get: \[z^2=-8(z^2-64)\] Can you then expand the brackets on the right hand side and tell me what you get?

OpenStudy (anonymous):

muliply 64 by 8

OpenStudy (anonymous):

Yes, don't forget to multiply \(z^2\) by -8 as well. You will get: \[z^2=-8z^2+512\] Now we can combine all \(z^2\) terms on the left hand side to get: \[9z^2=512\] Then divide by 9 to get: \[z^2=\frac{512}{9}\] And hence \(z=\sqrt{\frac{512}{9}}\) which if you simplify is \(z=\frac{16\sqrt{2}}{3} \) Do you understand?

OpenStudy (anonymous):

When I got it I got 512/9 and put it in and they said it was wrong.

OpenStudy (phi):

did you take the square root btw, you get ± (two answers)

OpenStudy (anonymous):

I cannot read the one where it was simplify

OpenStudy (anonymous):

\[z=\pm \sqrt{\frac{512}{9}}=\pm\frac{16\sqrt{2}}{3} \] The last part is the bit after the simplify.

OpenStudy (anonymous):

thanks I will try it